bella2025:

Write an equation in slope-intercept form of the line that passes through (-5,\ 19 ) and (5,\ 13)

MusicGeek:

Slope-intercept form is y=mx+b. To find m you would subtract the two points. 19-13 over -5-5. When you do that you should get 6/-10. Which you could reduce to -3/5. When you have that, use one of the points and plug them in for y and x to find b. Both points should get u the same answer of b=16. So you just have to put the numbers -3/5 and 16 in y=mx+b format. Which would be y=-3/5+16.

bella2025:

@musicgeek wrote:
Slope-intercept form is y=mx+b. To find m you would subtract the two points. 19-13 over -5-5. When you do that you should get 6/-10. Which you could reduce to -3/5. When you have that, use one of the points and plug them in for y and x to find b. Both points should get u the same answer of b=16. So you just have to put the numbers -3/5 and 16 in y=mx+b format. Which would be y=-3/5+16.
thanks

MusicGeek:

@bella2025 wrote:
@musicgeek wrote:
Slope-intercept form is y=mx+b. To find m you would subtract the two points. 19-13 over -5-5. When you do that you should get 6/-10. Which you could reduce to -3/5. When you have that, use one of the points and plug them in for y and x to find b. Both points should get u the same answer of b=16. So you just have to put the numbers -3/5 and 16 in y=mx+b format. Which would be y=-3/5+16.
thanks
You're welcome